### v1

# Resolving Quantum Criticality in the Honeycomb Hubbard Model

Resolving Quantum Criticality in the Honeycomb Hubbard Model
Fo-Hong Wang Tsung-Dao Lee Institute, Key Laboratory of Artificial Structures and Quantum Control (Ministry of Education), School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China    Fanjie Sun Tsung-Dao Lee Institute, Key Laboratory of Artificial Structures and Quantum Control (Ministry of Education), School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China    Chenghao He Tsung-Dao Lee Institute, Key Laboratory of Artificial Structures and Quantum Control (Ministry of Education), School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China    Xiao Yan Xu xiaoyanxu@sjtu.edu.cn Tsung-Dao Lee Institute, Key Laboratory of Artificial Structures and Quantum Control (Ministry of Education), School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China Hefei National Laboratory, Hefei 230088, China
The interplay between Dirac fermions and electronic correlations on the honeycomb lattice hosts a fundamental quantum phase transition from a semimetal to a Mott insulator, governed by the Gross-Neveu-Heisenberg (GNH) universality class. Despite its importance, consensus on the precise critical exponents remains elusive due to severe finite-size effects in numerical simulations and the lack of conformal bootstrap benchmarks. Here we try to resolve this long-standing controversy by performing projector determinant quantum Monte Carlo (QMC) simulations on lattices of unprecedented size, reaching 10,368 sites. By developing a novel projected submatrix update algorithm, we achieve a significant algorithmic speedup that enables us to access the thermodynamic limit with high precision. We observe that the fermion anomalous dimension and the correlation length exponent converge rapidly, while the boson anomalous dimension exhibits a systematic size dependence that we resolve via linear extrapolation. To validate our analysis, we perform parallel large-scale simulations of the spinless $t$-$V$ model on the honeycomb lattice, which belongs to the Gross-Neveu-Ising class. Our results for the $t$-$V$ model, including the first QMC determination of the fermion anomalous dimension, show agreement with conformal bootstrap predictions, thereby corroborating the robustness of our methodology. Our work provides state-of-the-art critical exponents for the honeycomb Hubbard model and establishes a systematic finite-size scaling workflow applicable to a broad class of strongly correlated quantum systems, paving the way for resolving other challenging fermionic quantum critical phenomena.
I Introduction
Understanding how quantum materials transition between distinct phases of matter is a central challenge in condensed matter physics. When massless Dirac fermions–the relativistic quasiparticles underlying graphene’s electronic properties and emerging in diverse quantum materials from $d$-wave superconductors to surface of topological insulators [1, 2, 3]–interact strongly, they can undergo quantum phase transitions that lie beyond conventional theoretical paradigms. A prototypical example is the interaction-driven semimetal-to-Mott-insulator transition in the honeycomb Hubbard model [4, 5, 6, 7, 8, 9], where increasing electron-electron repulsion spontaneously generates mass and magnetic order. This transition belongs to the Gross-Neveu-Heisenberg (GNH) universality class [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and connects fundamental questions in quantum materials to high-energy physics concepts such as chiral symmetry breaking and dynamical mass generation.
Despite decades of effort, consensus on the precise critical exponents governing this transition remains elusive, as illustrated in Fig. 1. Analytical approaches–including $\epsilon$ expansions, large-$N$ expansions, and functional renormalization group (FRG) calculations–yield estimates that are only partially compatible, with a particularly pronounced spread in the fermion anomalous dimension $\eta_{\psi}$ [10, 11, 12, 13]. Quantum Monte Carlo (QMC) results are even less uniform: for the honeycomb Hubbard model, most studies report a boson anomalous dimension $\eta_{\phi}\lesssim 0.65$ [14, 15, 16, 17, 18]–approximately 40% smaller than many analytical predictions [10, 11, 12, 13]–while different lattice realizations in the same universality class yield noticeably larger values. In contrast, for the $N=4$ Gross-Neveu-Ising (GNI) universality class [26, 27, 28, 29, 30, 31, 15, 32, 33, 34, 35, 36], large-scale QMC simulations [36] have produced critical exponents consistent with conformal bootstrap predictions [31], highlighting the achievable precision when finite-size effects are properly controlled.
Figure 1: Summary of critical exponents $\nu$, $\eta_{\phi}$, and $\eta_{\psi}$ reported in the literature and obtained in this work. a, $N=8$ Gross-Neveu-Heisenberg universality class [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. b, $N=4$ Gross-Neveu-Ising universality class [26, 27, 28, 29, 30, 31, 15, 32, 33, 34, 35, 36]. The gray horizontal dashed lines divide the plot into three blocks: analytical results (top), QMC results for the same model as in this work (middle), and QMC results for other models in the same universality class (bottom). CB denotes conformal bootstrap [31]. See Supplemental Material Tables S5 and S6 for the numerical values and methodological details.
The root cause of the GNH controversy lies in severe finite-size effects that render extracted exponents highly sensitive to the choice of system sizes and fitting procedures. The small leading-correction-to-scaling exponent $\omega\approx 0.3$ arises from irrelevant operators associated with competing interaction channels [11] and non-universal analytic contributions grounded in the large $\eta_{\phi}$ [19], which also induce large statistical fluctuations in susceptibilities sampled from QMC [22]. Alternative simulation setups have been proposed to mitigate these effects, including extended-hopping models that enlarge the linear dispersion region [37, 38, 39, 20] and elective-momentum ultrasize QMC methods [40]. Remarkably, SLAC-fermion formulations produce high-precision exponents and suggest that honeycomb estimates drift toward SLAC values once the smallest system sizes are excluded [20], though a noticeable discrepancy with analytical predictions persists. However, the honeycomb Hubbard model remains the minimal model for the GNH transition with direct experimental relevance to graphene and related two-dimensional materials, making it essential to resolve the controversy within this original setting. While hybrid Monte Carlo (HMC) simulations [16] have achieved lattices with $L>100$, ergodicity issues restrict reliable results to high temperatures, preventing direct access to ground-state properties. Their extracted correlation length exponent $\nu$ is consistent with theoretical predictions, but $\eta_{\phi}$ still deviates significantly from analytical estimates, and $\eta_{\psi}$ remains undetermined [16].
In this work, we revisit this long-standing problem with the numerically exact projector determinant quantum Monte Carlo (PQMC) [41, 42, 43] method on lattices up to $72\times 72\times 2$ sites—twice the linear dimension of previous PQMC study [18]. This unprecedented scale is enabled by a new submatrix update algorithm that optimizes cache utilization in CPU operations [44, 45, 46], achieving state-of-the-art performance without introducing any additional approximations. Through a straightforward yet systematic finite-size scaling (FSS) analysis, we unveil qualitatively distinct convergence patterns for the three critical exponents and extract their definitive values accordingly. Parallel simulations of the spinless $t$-$V$ model belonging to $N=4$ GNI universality class validate our approach and demonstrate that scaling up to such large sizes reveals the characteristic finite-size behaviors essential for resolving fermionic quantum criticality.
II Results
II.1 Semimetal-Mott-insulator transition in the honeycomb Hubbard model
We study the half-filled Hubbard model on the honeycomb lattice. The Hamiltonian is given by $\hat{H}=-t\sum_{\langle i,j\rangle,\sigma}(c_{i\sigma}^{\dagger}c_{j\sigma}+\mathrm{H.c.})+U\sum_{i}\left(\hat{n}_{i\uparrow}-\frac{1}{2}\right)\left(\hat{n}_{i\downarrow}-\frac{1}{2}\right)$, where $t$ is the nearest-neighbor hopping amplitude and $U$ is the Hubbard interaction strength. We impose periodic boundary conditions on an $L\times L$ honeycomb lattice with two sublattices, yielding $N_{\mathrm{s}}=2L^{2}$ lattice sites. At half filling, the low-energy excitations are spin-1/2 interacting Dirac fermions at the Dirac points $\mathrm{K}$ and $\mathrm{K^{\prime}}$. Increasing $U$ drives a continuous Mott transition belonging to the $N=8$ GNH universality class, from a semimetal to an antiferromagnetic (AFM) insulator with spontaneous spin-rotation symmetry breaking [4, 6, 7, 8, 9, 14, 15, 16, 17, 18, 19, 20], as illustrated in Fig. 2a. We employ the sign-problem-free PQMC [41, 42, 43] to obtain the ground-state properties of this model, with the electron number fixed to $N_{\mathrm{s}}$ to enforce half filling. The formulation, optimization, implementation, and parameter choices of the algorithm are provided in Methods and Supplemental Material [47].
To detect AFM long-range order on finite lattices, we define the correlation function in momentum space:
$C(\mathbf{k})=\frac{1}{L^{2}}\sum_{\mathbf{r}}e^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\left\langle\mathbf{\hat{O}}\left(\mathbf{r}\right)\cdot\mathbf{\hat{O}}\left(0\right)\right\rangle,$ (1)
where $\mathbf{r}$ runs over all unit cells and $\mathbf{\hat{O}}(\mathbf{r})=[\mathbf{\hat{S}}_{A}(\mathbf{r})-\mathbf{\hat{S}}_{B}(\mathbf{r})]/2$ is the local vector order parameter for AFM order. Here, we define the local spin operator as $\mathbf{\hat{S}}_{\alpha}(\mathbf{r})=\frac{1}{2}c_{\mathbf{r}\alpha\mu}^{\dagger}\bm{\sigma}_{\mu\nu}c_{\mathbf{r}\alpha\nu}$, where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, $\mu=\uparrow,\downarrow$, and $\alpha\in\{A,B\}$. At $\mathbf{k}=0$, the correlation function provides a finite-size proxy of the order parameter,
$m_{\mathrm{AFM}}^{2}\equiv\frac{1}{N_{\mathrm{s}}^{2}}\left\langle\left(\sum_{i\in A}\mathbf{\hat{S}}_{i}-\sum_{i\in B}\mathbf{\hat{S}}_{i}\right)^{2}\right\rangle=C(\mathbf{k}=0).$ (2)
A finite squared staggered magnetization $m^{2}_{\mathrm{AFM}}$ in the thermodynamic limit (TDL) indicates long-range AFM order, whereas it vanishes in the disordered phase.
Figure 2: Phase diagram and crossing-point analysis of the correlation ratio. a, Schematic phase diagram of the interaction-driven semimetal-Mott insulator transition with a Gross-Neveu-Heisenberg critical point. b, Correlation ratio $R_{\mathrm{AFM}}^{(1,0)}$ as a function of $U$ for various system sizes $L$. Horizontal black error bars mark the crossing points $U^{\times}$ of size pairs $(L,2L)$, determined by quadratic interpolation of the data for each size. Inset: power-law extrapolation of the crossing points.
Around the critical point, we employ the correlation ratio, an RG-invariant quantity constructed from the spin-spin correlation function, to sensitively probe the phase transition [19, 48, 20]:
$R_{\mathrm{AFM}}^{(n_{1},n_{2})}=1-\frac{C(\frac{n_{1}\mathbf{b}_{1}}{L}+\frac{n_{2}\mathbf{b}_{2}}{L})}{C(0)},$ (3)
where $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are the reciprocal lattice vectors. The integers $n_{1}$ and $n_{2}$ are typically chosen to be small (e.g., $|n_{1,2}|\leq 2$) and we will focus on $(n_{1},n_{2})=(1,0)$. In the TDL, $R_{\mathrm{AFM}}$ vanishes in the disordered phase and approaches unity in the ordered phase. At the critical point, $R_{\mathrm{AFM}}$ converges to a universal constant $R_{\mathrm{AFM}}^{*}$, signaled by the crossing of $R_{\mathrm{AFM}}$-$U$ curves for various $L$.
For finite systems, the crossing point between a pair of system sizes exhibits a drift due to scaling corrections [49, 19]. As shown in Fig. 2b, only for $L\gtrsim 36$ do the $R_{\mathrm{AFM}}$-$U$ curves approximately cross at a common point, indicating strong finite-size effects in the honeycomb Hubbard model. The final crossing point ($\sim 3.74$) is consistent with the transition scale estimated from the decay of the real-space spin-spin correlation function at maximum distance (see Supplemental Material [47] Section S6). To quantify this drift, we denote the crossing point between sizes $L$ and $2L$ as $U^{\times}(L,2L)$ (horizontal bars in Fig. 2b) and fit $U^{\times}(L,2L)=U_{c}^{\times}+a(2L)^{-b}$. The inset of Fig. 2b shows that this power-law extrapolation yields $U_{c}^{\times}=3.693(7)$ and $b_{\mathrm{fit}}=1.26(5)$. For an RG-invariant quantity, the drift exponent is expected to scale as $b=1/\nu+\min(\omega,2-z-\eta_{\phi})$ [49], where $\omega$ is the correction-to-scaling exponent from irrelevant operators and $z=1$ for the Lorentz-invariant Gross-Neveu criticality. Using the data-collapse estimates $\nu=1.11(9)$ and $\eta_{\phi}=0.715(15)$ obtained later with the data from the largest sizes, and noting that one-loop $\epsilon$-expansion studies suggest $\omega\sim 0.3$ for the $N=8$ GNH fixed point [11], we obtain $b_{\mathrm{collapse}}=1.19(5)$, consistent with $b_{\mathrm{fit}}$ within error bars.
II.2 Gross-Neveu-Heisenberg criticality
Figure 3: Sliding-window data-collapse analysis for the Hubbard model. Columns show three representative fit windows with $L_{\mathrm{max}}=36$, 54, and 72 (see Supplemental Material [47] Section S6 for all windows). The $L$-axis above each panel indicates the sizes included in the fit, with greyed-out sizes excluded. Rows: $R_{\mathrm{AFM}}^{(1,0)}$ collapse yielding $(U_{c},\nu)$ (top), $m^{2}_{\mathrm{AFM}}$ collapse yielding $\eta_{\phi}$ (middle), and $\langle|G_{AB}|\rangle$ collapse yielding $\eta_{\psi}$ (bottom).
With the crossing-point analysis providing a rough initial estimate of $U_{c}$, we now extract the GNH critical exponents via a systematic FSS analysis. The critical point $U_{c}$ and exponent $\nu$ are obtained from the ansatz for the RG-invariant correlation ratio:
$R_{\text{AFM}}^{(n_{1},n_{2})}(U,L)$ $=f^{R}\left(uL^{1/\nu},L^{-\omega}\right)$ (4)
$\approx f_{0}^{R}\left(uL^{1/\nu}\right)+L^{-\omega}f_{1}^{R}\left(uL^{1/\nu}\right),$
where $u=U-U_{c}$ is the distance from criticality, and $\omega$ denotes the leading correction-to-scaling exponent arising from irrelevant operators [11] or non-universal analytic contributions [19] (see also the preceding discussion). The second line expands $f^{R}$ to first order in $L^{-\omega}$. In our analysis, only $f_{0}^{R}$ is fitted while the correction term is neglected—a procedure termed data collapse, as data at different $U$ and $L$ collapse onto a single curve when $U_{c}$ and $\nu$ are properly chosen. We find that including the $L^{-\omega}$ correction leads to unstable fits with large reduced chi-squared values, whereas the leading-order collapse already yields visually satisfactory agreement.
The anomalous dimensions are extracted from the correlation functions. For the boson anomalous dimension $\eta_{\phi}$, the squared staggered magnetization obeys the FSS ansatz
$m^{2}_{\mathrm{AFM}}\left(R_{\text{AFM}}^{(n_{1},n_{2})},L\right)=L^{-1-\eta_{\phi}}f_{0}^{m}\left(R_{\text{AFM}}^{(n_{1},n_{2})}\right).$ (5)
Instead of fitting against $uL^{1/\nu}$, the scaling is analyzed with respect to $R_{\text{AFM}}^{(n_{1},n_{2})}$ [19, 48, 20], which bypasses prior knowledge of $U_{c}$ and $\nu$ and yields a more stable single-parameter fit. The fermion anomalous dimension $\eta_{\psi}$ is extracted from the off-diagonal Green’s function in momentum space, $G_{AB}(\mathbf{k})=\sum_{\mathbf{r}}e^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\langle c_{\mathbf{r}A\mu}^{\dagger}c_{0B\mu}\rangle$, where $\mu=\uparrow$ or $\downarrow$. The averaged modulus $\left\langle\left|G_{AB}(\mathrm{K}+\delta\mathbf{k})\right|\right\rangle_{\delta\mathbf{k}}$ over the six momenta nearest to the Dirac point $\mathrm{K}$ follows the FSS ansatz [37, 38, 20, 14, 18]
$\left\langle\left|G_{AB}(\mathrm{K}+\delta\mathbf{k})\right|\right\rangle_{\delta\mathbf{k}}\left(R_{\text{AFM}}^{(n_{1},n_{2})},L\right)=L^{-\eta_{\psi}}f_{0}^{G}\left(R_{\text{AFM}}^{(n_{1},n_{2})}\right).$ (6)
We provide an argument for this ansatz, including the rationale for taking the modulus, in Supplemental Material [47] Section S5.
Figure 4: Fitting-window dependence of critical parameters for the Hubbard and spinless $t$-$V$ models. ad, Hubbard model; eh, $t$-$V$ model. Each point represents the fitted value of $U_{c}$ (or $V_{c}$), $\nu$, $\eta_{\phi}$, or $\eta_{\psi}$ obtained from a fitting window of six consecutive sizes, with $L_{\mathrm{max}}$ denoting the largest size in each window. Horizontal lines and shaded bands indicate final estimates and confidence intervals; dashed lines show linear extrapolations to $1/L_{\mathrm{max}}\to 0$.
Given the extended range of accessible system sizes, selecting an appropriate fitting window is not obvious a priori. A common strategy in previous QMC studies is to successively discard the smallest sizes until $\chi^{2}_{\mathrm{red}}$ (see Methods) becomes acceptable [19, 23, 48, 20]. We refer to this practice as a shrinking-window analysis: the window width (i.e. the number of consecutive system sizes included in the fit) decreases while the largest size $L_{\max}$ is fixed. Notably, Ref. [20] reported a monotonic drift of $\eta_{\phi}$ with increasing the smallest size $L_{\min}$, trending toward the SLAC-fermion estimate. To learn this window dependence systematically, we scan over both upper and lower cutoffs $L_{\min}$ and $L_{\max}$, finding that overly narrow windows usually yield larger error bars (since data are sparser and thus impose fewer constraints on the fit), whereas overly wide windows tend to increase $\chi^{2}_{\mathrm{red}}$ (see also the tables in Refs. [19, 48]). The complete scan over fitting windows is shown in the heatmaps in Supplemental Material [47] Section S9.
Based on the above observations, we adopt a sliding-window data-collapse analysis: we fix the window width at six consecutive sizes and slide it across all accessible $L$ (choosing a window width of five yields similar results). Within each window, we fit the scaling forms in Eqs. (4)–(6) and extract $(U_{c},\nu,\eta_{\phi},\eta_{\psi})$ following Methods, labeling each window by its largest size $L_{\mathrm{max}}$. Figure 3 illustrates this procedure for the Hubbard model: as the window slides from left to right toward larger $L$, the fitted $U_{c}$, $\eta_{\phi}$, and $\eta_{\psi}$ drift monotonically, whereas $\nu$ appears to oscillate. For the $L_{\mathrm{max}}=72$ window (rightmost column), we obtain $U_{c}=3.7422(34)$, $\nu=1.11(5)$, $\eta_{\phi}=0.715(15)$, and $\eta_{\psi}=0.1881(33)$. We also perform the same FSS analysis for the spinless $t$-$V$ model on the honeycomb lattice, whose critical point belongs to the $N=4$ GNI universality class, with system sizes up to $L=48$ (see Supplemental Material [47] Section S7 for details).
To quantify the trends in the fitted parameters systematically, we plot them against $1/L_{\mathrm{max}}$ for both models in Fig. 4. Three distinct patterns emerge as $1/L_{\mathrm{max}}$ decreases: rapid convergence to a plateau ($V_{c}$ in the $t$-$V$ model, $\eta_{\psi}$ in the Hubbard model), stable oscillations ($\nu$ in both models, $\eta_{\psi}$ in the $t$-$V$ model), and persistent monotonic drift ($U_{c}$ in the Hubbard model, $\eta_{\phi}$ in both models). For the first two patterns, we estimate definitive values by averaging over converged windows, with uncertainties taken conservatively from the observed spread. For the last kind of pattern, we extrapolate linearly to $1/L_{\mathrm{max}}\to 0$. This procedure determines the final estimation of all critical parameters in this work: $U_{c}=3.664(5)$, $\nu=1.11(9)$, $\eta_{\phi}=0.79(2)$, and $\eta_{\psi}=0.1888(40)$ for the Hubbard model, and $V_{c}=1.3378(11)$, $\nu=0.87(6)$, $\eta_{\phi}=0.49(3)$, and $\eta_{\psi}=0.0905(35)$ for the $t$-$V$ model.
III Discussion
By combining unprecedented system sizes with a systematic sliding-window scaling analysis, we determine critical exponents with well-controlled uncertainties for quantum criticality in the honeycomb Hubbard model. A central finding is that different observables approach the TDL at markedly different rates. For the two-parameter correlation-ratio collapse, the best-fit $U_{c}$ drifts appreciably across sliding windows (Figs. 3 and 4), similar to the drifting crossing points in Fig. 2. Meanwhile, $\nu$ remains stable across sliding windows, yielding a comparatively wide confidence interval. Turning to the anomalous dimensions, $\eta_{\psi}$ rapidly plateaus within our numerical precision, whereas $\eta_{\phi}$ exhibits a monotonic drift. This difference is attributed to stronger finite-size corrections in the squared staggered magnetization, arising from the much larger boson anomalous dimension [49, 19]. Specifically, the rapidly decaying correlations $\langle\mathbf{\hat{O}}(\mathbf{r})\cdot\mathbf{\hat{O}}(0)\rangle$ induce a relatively large background term $B(U)$ in the structure factor $S_{\text{AFM}}\equiv N_{\mathrm{s}}m^{2}_{\mathrm{AFM}}$, arising from non-universal short-range correlations. At the critical point, one can write $S_{\text{AFM}}=L^{1-\eta_{\phi}}\left[A+B(U_{c})L^{-\omega_{\eta}}\right]$, with $\omega_{\eta}\equiv 1-\eta_{\phi}$ an effective correction-to-scaling exponent.
Placed in the broader context of prior studies, our results are compared with representative analytical calculations and large-scale QMC simulations in Fig. 1. For the $N=4$ GNI transition realized in the spinless $t$-$V$ model, our $\nu$ and $\eta_{\phi}$ agree with both conformal bootstrap (CB) predictions [31] and fermion-bag QMC simulations [36]. Most notably, we provide the first QMC estimate of $\eta_{\psi}$ in this model, which is consistent with the CB prediction within uncertainties. For the $N=8$ GNH transition in the honeycomb Hubbard model, the wide confidence interval of our $\nu$ encompasses most prior QMC and analytical values. Our $\eta_{\phi}$ moves closer to analytical predictions and aligns with most other lattice realizations in the same universality class [25, 24, 23, 22, 20], and our $\eta_{\psi}\approx 0.19$ remains consistent with earlier QMC values near $0.2$, though a gap to analytical predictions persists for both anomalous dimensions.
The sliding-window analysis in Fig. 4 offers internal validation of our final estimates, by tracking how the fitted parameters evolve as system size increases. When we restrict the Hubbard data to $L\leq 36$, from Figs. 3a, d and g, we obtain $U_{c}=3.8187(30)$, $\nu=1.058(27)$, $\eta_{\phi}=0.628(7)$, and $\eta_{\psi}=0.1948(15)$, values closer to earlier PQMC studies on the same model [18, 20]. As the sliding window moves to larger sizes, $\eta_{\phi}$ increases monotonically, and $\eta_{\psi}$ decreases slightly, all shifting from earlier honeycomb Hubbard estimates toward analytical predictions. This systematic drift demonstrates that finite-size effects are the primary source of discrepancies between earlier honeycomb Hubbard results and other estimates. More broadly, beyond serving as an internal consistency check in the present work, the sliding-window analysis offers a transparent way to assess the stability of FSS estimates as system size grows. Compared with the conventional shrinking-window strategy, where $L_{\max}$ is fixed and $L_{\min}$ is thereby bounded from above, the sliding-window shifts the entire fit interval toward larger sizes and thus provides a more natural extrapolation to the TDL.
In summary, the combination of our algorithmic advances and systematic scaling analysis directly addresses a long-standing bottleneck in fermionic quantum criticality: the limited accessible system sizes that prevent reliable extraction of universal critical exponents. By pushing lattice simulations to unprecedented scales–10,368 sites for the honeycomb Hubbard model–we demonstrate that many apparent controversies in the literature stem from insufficient system sizes rather than fundamental methodological differences. This capability is particularly valuable for fermionic quantum critical points where severe finite-size effects have hindered consensus. Our sliding-window analysis further reveals that finite-size corrections vary strongly among different observables (Fig. 4)–a diagnostic insight that can guide future studies in identifying which quantities require extrapolation versus direct averaging.
Looking ahead, the sliding-window protocol developed here can prove broadly useful for extracting critical exponents in other quantum phase transitions, especially when finite-size effects are severe. With both the computational power to reach larger sizes and a robust protocol to extract asymptotic exponents, the path toward resolving other outstanding problems in strongly correlated quantum systems–particularly those involving competing orders, emergent gauge structures, or unconventional quantum criticality–is now significantly clearer.
IV Methods
IV.1 Projector determinant quantum Monte Carlo
The projector determinant quantum Monte Carlo (PQMC) algorithm evaluates ground-state properties by imaginary-time projection of a trial Slater determinant $\left|\Psi_{T}\right\rangle$ [41, 42, 43]:
$\langle{\hat{O}}\rangle=\frac{\left\langle\Psi_{T}\left|e^{-\Theta\hat{H}}\hat{O}e^{-\Theta\hat{H}}\right|\Psi_{T}\right\rangle}{\left\langle\Psi_{T}\left|e^{-2\Theta\hat{H}}\right|\Psi_{T}\right\rangle}.$ (7)
We use the symmetric Trotter decomposition to discretize the projection length $\beta\equiv 2\Theta=L_{\tau}\Delta_{\tau}$: $e^{-\Delta_{\tau}\hat{H}}\approx e^{-\Delta_{\tau}\hat{H}_{0}/2}e^{-\Delta_{\tau}\hat{H}_{I}}e^{-\Delta_{\tau}\hat{H}_{0}/2}$. The trial wave function $|\Psi_{T}\rangle$ is chosen as the ground state of the non-interacting Hamiltonian $\hat{H}_{0}$, with small random hoppings added on each nearest-neighbor bond to lift energy level degeneracies at Dirac points. The interaction term $\hat{H}_{I}$ is decoupled via the Hubbard-Stratonovich (HS) transformation. For the repulsive Hubbard model, we employ the discrete HS transformation in the density channel [50]:
$e^{-\Delta_{\tau}U\left(\hat{n}_{i\uparrow}-\frac{1}{2}\right)\left(\hat{n}_{i\downarrow}-\frac{1}{2}\right)}={\gamma}\sum_{s_{i}=\pm 1}e^{\mathrm{i}{\alpha}s_{i}\left(\hat{n}_{i\uparrow}+\hat{n}_{i\downarrow}-1\right)},$ (8)
where ${\gamma}=\frac{1}{2}e^{\Delta_{\tau}U/4}$ and $\cos{\alpha}=e^{-\Delta_{\tau}U/2}$. For the spinless $t$-$V$ model, we decouple the nearest-neighbor repulsion to the hopping channel [51, 52]:
$e^{-\Delta\tau V\left(\hat{n}_{j}-\frac{1}{2}\right)\left(\hat{n}_{k}-\frac{1}{2}\right)}=\zeta\sum_{s_{jk}=\pm 1}e^{\lambda s_{jk}\left(c_{j}^{\dagger}c_{k}+c_{k}^{\dagger}c_{j}\right)},$ (9)
where $\zeta=\frac{1}{2}e^{-\frac{V\Delta\tau}{4}}$ and $\cosh\lambda=e^{\frac{V\Delta\tau}{2}}$. In the Hubbard model, the auxiliary fields are defined on lattice sites, whereas in the $t$-$V$ model, they reside in nearest-neighbor bonds. We denote by $\mathbf{s}$ the complete set of auxiliary fields across all spatial locations and imaginary-time slices.
Systematic errors in PQMC arise primarily from the Trotter discretization $\Delta_{\tau}$ and the finite projection length $\beta$. We control these errors by performing convergence tests for both models (see Supplemental Material [47] Section S4 for details). For the Hubbard model, we adopt $\Delta_{\tau}t=0.1$ [19, 18, 20], whereas for the $t$-$V$ model, a smaller time step $\Delta_{\tau}t=0.05$ is required [15]. The projection length is scaled with the linear system size as $\beta t=L+12$ for both models, which is sufficient to render systematic errors smaller than statistical fluctuations.
IV.2 Submatrix update algorithm for PQMC
We first briefly review the traditional fast update algorithm for $T\equiv(LR)^{-1}$. More details can be found in Supplemental Material [47] Section S1. After tracing out the fermionic degrees of freedom, the Monte Carlo weight for configuration $\mathbf{s}$ is proportional to a fermionic determinant $\det[L(\tau)R(\tau)]$, where $R(\tau)$ (size $N\times N_{p}$) and $L(\tau)$ (size $N_{p}\times N$) are rectangular matrices encoding the imaginary-time propagation from the trial state to time slice $\tau$ from 0 and $\beta$, respectively. Here $N$ typically denotes the total number of lattice sites and $N_{p}$ is the number of particles per spin block.
In PQMC, the most computationally intensive part is updating $T$ upon changes in auxiliary fields [44]. A local update of the auxiliary field at specific spatial position $x$ and time slice $\tau$ modifies $R$ by a rank-$k$ correction, $R^{\prime}=(I+\Delta)R$, and the Metropolis acceptance ratio involves the determinant ratio $r=\det[LR^{\prime}]/\det[LR]$. The Sherman-Morrison-Woodbury formula allows for a fast update of $T$:
$T^{\prime}=T-TU(I_{k}+VU)^{-1}V,$ (10)
where $U=L\Delta P_{N\times k}$ and $V=P_{k\times N}RT$ are small rectangular matrices. Here, $\Delta$ is a sparse diagonal matrix with only $k$ non-zero entries at the specific lattice sites being updated ($x=\{x_{1},\ldots,x_{k}\}$), and we introduce the index matrices
$P_{N\times k}=\left[e_{x_{1}}|e_{x_{2}}|\cdots|e_{x_{k}}\right]\text{ and }P_{k\times N}=P_{N\times k}^{T},$ (11)
where $e_{x_{j}}=[0,\ldots,0,1,0,\ldots,0]^{T}$ is the $x_{j}$-th unit vector of length $N$ (several properties of the index matrices are discussed in Supplemental Material [47] Section S1). The acceptance ratio is given by $r=\det[I_{k}+VU]$. The per-move cost is $\mathcal{O}(N_{p}^{2})$ and the per-time-slice cost is $\mathcal{O}(N_{p}^{2}N)$.
While this fast update formula enables efficient per-move computation, it relies heavily on rank-$k$ matrix-vector operations (BLAS-1&2) that are not cache-friendly. To overcome this bottleneck, delayed update algorithms, accumulating multiple local updates before performing a single, cache-efficient, higher-rank batch operation (BLAS-3), were proposed [44, 45, 46]. In Supplemental Material [47] Section S2, we provide a self-contained review of delay-G [44], submatrix-G [45], and delay-T [46] algorithms.
In this work, we introduce a new “submatrix-T” update algorithm which outperforms all existing delayed update algorithms for PQMC simulations. As illustrated in Fig. 5, a series of accepted local updates are accumulated within a block (dubbed “delayed block”) before performing a full update. We consider how to write down the update formula (as in Eq. (10)) and ratio formula using the original $T$ without any updating. To express these formulas, it is convenient to introduce a superscript $\cdot]^{(i)}$ which labels quantities after $i$ accepted moves or quantities at the proposed move after i1i-1 accepted moves in the current delayed block. The lattice sites being updated or may be updated is denoted by $x^{(i)}=\{x_{1}^{(i)},\ldots,x_{k}^{(i)}\}$, which further determines the difference $\Delta^{(i)}$ and the index matrix $P_{N\times k}^{(i)}$. Other examples include the update object $T^{(i)}=(LR^{(i)})^{-1}$ and intermediate matrices like $U^{(i)}\equiv L^{(i-1)}\Delta^{(i)}P_{N\times k}^{(i)}$ and $V^{(i)}\equiv P_{k\times N}^{(i)}R^{(i-1)}T^{(i-1)}$.
Figure 5: Schematic illustration of the submatrix-T update algorithm. At a specific time slice, there are $N/k$ proposed local updates in total, which is divided into several delayed blocks of size $n_{d}$. Within each delayed block, we do not update $T$ (indicated by dashed arrows) but instead perform a recursive update of the submatrix $\Gamma^{-1}$ (indicated by solid arrows in black), which helps in calculating not only the intermediate acceptance ratios $r^{(i)}$ but also the final $T^{(n_{a})}$ (indicated by solid arrows in grey), where $n_{a}$ denotes the total accepted moves in the current delayed block.
The most essential insight for designing submatrix-T update is that $L$, $R$ and $T$ can find their general term formulas after arbitrary number of accepted local updates. Within our sweep strategy, at a given time slice, $L$ does not update so $L^{(i)}=L^{(0)}=L$ (c.f. Eq. (S14) in Supplemental Material [47] Section S1). Since different local updates operate on disjoint lattice sites, the matrix $R$ can be updated by successive multiplication:
$R^{(i)}=\prod_{m=1}^{i}(I+\Delta^{(m)})R^{(0)}=\left(I+\Delta^{(i)}_{\mathrm{all}}\right)R^{(0)},$ (12)
where $\Delta^{(i)}_{\mathrm{all}}=\sum_{m=1}^{i}\Delta^{(m)}$ is the accumulated diagonal matrix containing $ik$ non-zero elements. This operation is computationally inexpensive, requiring only element-wise scaling of the affected rows of $R^{(0)}$. Another result is that at next proposed move, since $P_{N\times k}^{(i+1)}$ do not overlap with $\Delta^{(i)}_{\mathrm{all}}$, we have $P_{N\times k}^{(i+1)}R^{(i)}=P_{N\times k}^{(i+1)}R^{(0)}$, or more generally,
$P_{N\times k}^{(j)}R^{(i)}=P_{N\times k}^{(j)}R^{(0)}\text{ provided that }j>i.$ (13)
The general term formula for $T$ directly follows from Eq. (12). After utilizing the Woodbury matrix identity, we obtain:
$T^{(i)}$ $=\big(L(I+\Delta^{(i)}_{\mathrm{all}})R^{(0)}\big)^{-1}$ (14)
$=T^{(0)}-T^{(0)}U_{N_{p}\times ik}^{(i)}$
$\quad\times\big(I_{ik}+V_{ik\times N_{p}}^{(i)}U_{N_{p}\times ik}^{(i)}\big)^{-1}V_{ik\times N_{p}}^{(i)}\,$,
where the accumulated intermediate matrices are defined as $U_{N_{p}\times ik}^{(i)}=L\Delta^{(i)}_{\mathrm{all}}P_{N\times ik}^{(i)}$ and $V_{ik\times N_{p}}^{(i)}=P_{ik\times N}^{(i)}R^{(0)}T^{(0)}$. Here, $P_{N\times ik}^{(i)}$ and $P_{ik\times N}^{(i)}$ are the accumulated index matrices which are constructed by concatenating the index matrices from each accepted update:
$P_{N\times ik}^{(i)}$ $=\left[P_{N\times k}^{(1)}|P_{N\times k}^{(2)}|\cdots|P_{N\times k}^{(i)}\right]$ (15)
$=\left[e_{x_{1}^{(1)}}|\cdots|e_{x_{k}^{(1)}}|e_{x_{1}^{(2)}}|\cdots|e_{x_{k}^{(2)}}|\cdots|e_{x_{1}^{(i)}}|\cdots|e_{x_{k}^{(i)}}\right],$
$P_{ik\times N}^{(i)}$ $=(P_{N\times ik}^{(i)})^{T}.$
Crucially, the order of columns in $P_{N\times ik}^{(i)}$ is determined by the sequence of accepted moves, and does not necessarily correspond to the ordered indices of the non-zero elements in $\Delta^{(i)}_{\mathrm{all}}$. This is permissible due to the permutation invariance: for any $k\times k$ permutation matrix $\mathbf{S}_{k}$, we have $\Delta=\Delta\left(P_{N\times k}\mathbf{S}_{k}\right)\left(\mathbf{S}_{k}^{T}P_{k\times N}\right)$.
The next step of designing the submatrix-T update algorithm is to figure out how to compute the acceptance ratio for next proposed move recursively. We start with the fast update ratio formula $r=\det[I_{k}+VU]$:
$r^{(i+1)}$ $=\frac{\det[LR^{(i+1)}]}{\det[LR^{(i)}]}=\det[I_{k}+V_{k\times N_{p}}^{(i+1)}U_{N_{p}\times k}^{(i+1)}]$ (16)
$=\det[I_{k}+\mathcal{V}_{k\times k}^{(i+1)}D_{k\times k}^{(i+1)}]\,$,
where we further factorize the $k\times k$ determinant with matrices $\mathcal{V}_{k\times k}^{(i+1)}\equiv P_{k\times N}^{(i+1)}F^{(i)}P_{N\times k}^{(i+1)}$ and $D_{k\times k}^{(i+1)}\equiv P_{k\times N}^{(i+1)}\Delta^{(i+1)}P_{N\times k}^{(i+1)}$. The $N\times N$ matrix $F^{(i)}\equiv R^{(i)}T^{(i)}L$ has a similar role as the Green’s function $G^{(i)}$ (which equals to $(I-F^{(i)})$), and it is hired as the update object in the submatrix-G algorithm (see Supplemental Material [47] Section S2 for detailed formulations). Its general term formula follows from Eqs. (12) and (14):
$F^{(i)}=$ $\left(I+\Delta_{\mathrm{all}}^{(i)}\right)F^{(0)}\Big[I-\Delta_{\mathrm{all}}^{(i)}P_{N\times ik}^{(i)}$ (17)
$\times\left(I_{ik}+P_{ik\times N}^{(i)}F^{(0)}\Delta_{\mathrm{all}}^{(i)}P_{N\times ik}^{(i)}\right)^{-1}P_{ik\times N}^{(i)}F^{(0)}\Big].$
Substituting this into $\mathcal{V}^{(i+1)}$ and utilizing Eq. (13) yields
$\mathcal{V}_{k\times k}^{(i+1)}$ $=P_{k\times N}^{(i+1)}F^{(i)}P_{N\times k}^{(i+1)}$ (18)
$=P_{k\times N}^{(i+1)}F^{(0)}P_{N\times k}^{(i+1)}$
$\quad-P_{k\times N}^{(i+1)}F^{(0)}P_{N\times ik}^{(i)}\big(\Gamma^{(i)}\big)^{-1}P_{ik\times N}^{(i)}F^{(0)}P_{N\times k}^{(i+1)}.$
where we identify the $ik\times ik$ submatrix $\Gamma^{(i)}$ defined as:
$\Gamma^{(i)}$ $=\left(P_{ik\times N}^{(i)}\Delta_{\mathrm{all}}^{(i)}P_{N\times ik}^{(i)}\right)^{-1}+P_{ik\times N}^{(i)}F^{(0)}P_{N\times ik}^{(i)}$ (19)
$=\left(\begin{array}[]{ccc}\left(D_{k\times k}^{(1)}\right)^{-1}\\ &\ddots\\ &&\left(D_{k\times k}^{(i)}\right)^{-1}\end{array}\right)+P_{ik\times N}^{(i)}F^{(0)}P_{N\times ik}^{(i)}.$
This submatrix $\Gamma^{(i)}$ is the same as that in the submatrix-G algorithm [45]. In addition to calculating acceptance ratios, it also aids in performing full update of $T$. Following Eq. (14) we have
$T^{(i)}=T^{(0)}-T^{(0)}\big(LP_{N\times ik}^{(i)}\big)\big(\Gamma^{(i)}\big)^{-1}\big(P_{ik\times N}^{(i)}R^{(0)}T^{(0)}\big).$ (20)
The first term of $\Gamma^{(i)}$ is a block-diagonal matrix determined by the difference $\Delta$ of each accepted move, while the second term simply comprises selected elements of $F^{(0)}$. As $i$ increases, $\Gamma^{(i)}$ simply expands and satisfies the following recursive relation:
$\Gamma^{(i+1)}=\begin{pmatrix}\Gamma^{(i)}&F_{ik,k}^{(i+1)}\\ F_{k,ik}^{(i+1)}&(D_{k\times k}^{(i+1)})^{-1}+F_{k,k}^{(i+1)}\end{pmatrix}$, (21)
with short-hand notation for blocks of $F^{(0)}$:
$F_{ik,k}^{(i+1)}$ $\equiv P_{ik\times N}^{(i)}F^{(0)}P_{N\times k}^{(i+1)},$ (22)
$F_{k,ik}^{(i+1)}$ $\equiv P_{k\times N}^{(i+1)}F^{(0)}P_{N\times ik}^{(i)},$
$F_{k,k}^{(i+1)}$ $\equiv P_{k\times N}^{(i+1)}F^{(0)}P_{N\times k}^{(i+1)}.$
Using the block matrix inversion identity, $(\Gamma^{(i)})^{-1}$ can also be updated efficiently [45]:
$\left(\Gamma^{(i+1)}\right)^{-1}=\begin{pmatrix}\left(\Gamma^{(i)}\right)^{-1}\Upsilon&-\left(\Gamma^{(i)}\right)^{-1}F_{ik,k}^{(i+1)}\Sigma\\ -\Sigma F_{k,ik}^{(i+1)}\left(\Gamma^{(i)}\right)^{-1}&\Sigma\end{pmatrix}$, (23)
where
$\Upsilon$ $\equiv\left(I+F_{ik,k}^{(i+1)}\Sigma F_{k,ik}^{(i+1)}\left(\Gamma^{(i)}\right)^{-1}\right),$ (24)
$\Sigma$ $\equiv\left(\left(D_{k\times k}^{(i+1)}\right)^{-1}+F_{k,k}^{(i+1)}-F_{k,ik}^{(i+1)}\left(\Gamma^{(i)}\right)^{-1}F_{ik,k}^{(i+1)}\right)^{-1}$
$=D_{k\times k}^{(i+1)}\left(I_{k}+\mathcal{V}_{k\times k}^{(i+1)}D_{k\times k}^{(i+1)}\right)^{-1}.$
We note that the matrix $(I_{k}+\mathcal{V}_{k\times k}^{i+1}D_{k\times k}^{(i+1)})$ also appears in the acceptance ratio in Eq. (16).
We use Eqs. (16) and (18) to compute acceptance ratios. We keep track of $(\Gamma^{(i)})^{-1}$ and use Eq. (23) to update it after each accepted move. Finally, we use Eqs. (12) and (20) to perform full update of $R$ and $T$ respectively after a full delayed block. These formulas form the core of the submatrix-T algorithm illustrated in Fig. 5.
Figure 6 benchmarks the update time of submatrix-T algorithm against fast update and other existing delayed update algorithms. Submatrix-T consistently outperforms all alternatives. For the $L=36$ honeycomb lattice in single-threaded execution, submatrix-T achieves more than $24\times$ speedup over fast update, and this speedup factor increases with system size, reaching higher acceleration for larger lattices. The superior performance stems from replacing cache-inefficient BLAS-1&2 operations with batched BLAS-3 routines, which not only improves cache utilization but also provides better amenability to parallel acceleration on multi-threaded (e.g., OpenMP) and GPU architectures. For implementation details including the workflow flowchart, computational complexity analysis, hardware specifications, and comprehensive $n_{d}$ optimization results, see Supplemental Material [47] Section S3. As a low-level optimization, our submatrix-T algorithm can be broadly applied to accelerate fermionic computations within the PQMC framework, including entanglement measures [53, 54, 55, 56], stabilizer Rényi entropy [57], and finite-temperature schemes evaluating Fock-state expectations [58, 59].
Figure 6: Per-sweep update time comparison for different update algorithms in PQMC simulations of the Hubbard model on honeycomb lattices. All delayed update algorithms of various system sizes use their respective optimal $n_{d}$ values, and the inset illustrates the speedups achieved by the delayed update algorithms compared to the fast update.
IV.3 Data collapse analysis
Finite-size scaling (FSS) characterizes critical behavior when the linear system size is comparable to the correlation length and thus serves as the most relevant length scale, i.e., $\xi\sim L$. One expects a singular quantity to scale as $Q(t,L)\sim L^{\sigma}g(tL^{1/\nu})$ with $t$ measuring the distance of coupling from criticality and $\sigma$ a universal exponent associated with $Q$. Within a fixed fitting window of system sizes and couplings around the transition, we perform a one-parameter data-collapse fit using the scaling ansatz
$Y(X,L)=L^{c_{2}}F\left[(X-X_{c})L^{c_{1}}\right],$ (25)
where the scaling function $F$ is a priori unknown but shared by all $L$. A successful data collapse is achieved when $Y/L^{c_{2}}$ for different $L$ falls onto a single curve when plotted against the scaling variable $(X-X_{c})L^{c_{1}}$.
The scaling form (25) is applied to the correlation ratio $R$, the squared order parameter $m^{2}_{\mathrm{AFM}}$, and the off-diagonal Green’s function $G_{AB}$. For the dimensionless $R$ [Eq. (4)], we set $X$ to the tuning parameter ($U$ for the Hubbard and $V$ for the $t$-$V$ model) and fix $c_{2}=0$ to extract $X_{c}$ and $c_{1}=1/\nu$. For $m^{2}_{\mathrm{AFM}}$ and $G_{AB}$ [Eqs. (5) and (6)], we use $X=R$ and fix $X_{c}=c_{1}=0$, yielding $Y(R,L)=L^{c_{2}}F(R)$. Here $c_{2}$ encodes the anomalous dimensions: $c_{2}=-(1+\eta_{\phi})$ for $m^{2}_{\mathrm{AFM}}$ and $c_{2}=-\eta_{\psi}$ for $G_{AB}$.
To avoid choosing an explicit ansatz for $F$, we adopt the Bayesian scaling analysis (BSA) based on Gaussian-process regression [60]. The scaling function $F$ is modeled as a Gaussian process defined by its kernel hyperparameters. The physical scaling parameters ($X_{c}$, $c_{1}$, $c_{2}$) and the hyperparameters are fitted jointly; in what follows, we refer to all of them collectively as the fitting parameters. We use the reduced chi-squared $\chi^{2}_{\mathrm{red}}=\chi^{2}/(N_{\mathrm{data}}-N_{\mathrm{free}})$ to evaluate the goodness of fit, where $N_{\mathrm{data}}$ is the number of data points in the fitting window and $N_{\mathrm{free}}$ is the number of unfixed fitting parameters. The chi-squared is computed as $\chi^{2}=\sum_{i}\left[F(X_{i})-Y_{i}\right]^{2}/E_{i}^{2}$, where $E_{i}^{2}=\sigma_{Y_{i}}^{2}+\left[F^{\prime}(X_{i})\right]^{2}\sigma_{X_{i}}^{2}-2F^{\prime}(X_{i})\sigma_{X_{i}Y_{i}}$ is the effective variance that accounts for uncertainties in both $X$ and $Y$ data, as well as their covariance $\sigma_{X_{i}Y_{i}}$ [61, 62]. We provide more details in the $\chi^{2}_{\mathrm{red}}$ calculation of the three kinds of collapses in Supplemental Material [47] Section S9.
We estimate the uncertainties of the fitting parameters via a bootstrap workflow, implemented as an open-source Julia package at https://github.com/wangfh5/BSAHelper, serving as a downstream utility for the BSA program.
  1. 1.
    Data resampling. For each data point, resample $Y$ by adding Gaussian noise according to its Monte-Carlo error bar. When the horizontal axis is the correlation ratio ($X=R$), we additionally resample $X$ in the same way.
  2. 2.
    Initial-parameter jittering. Randomize the initial guesses of the fitting parameters (e.g., $X_{c}$, $c_{1}$, and $c_{2}$) by drawing each parameter uniformly from an interval $p_{0}-\Delta p,\,p_{0}+\Delta p]$.
  3. 3.
    Single BSA fit. Run BSA program once for the resampled dataset and record the best-fit values of the fitting parameters. The input errors of $Y$ are set to zero to avoid double counting the statistical uncertainty already injected by the resampling. We turn off the program’s internal uncertainty estimation for the fitting parameters.
  4. 4.
    Bootstrap statistics. Repeat steps 1–3 for $N_{\mathrm{boot}}$ trials and keep successful fits. For each fitting parameter, compute the bootstrap averages and use the standard deviations as the uncertainties.
  5. 5.
    Reconstruction and chi-squared test. Using the original Monte-Carlo-averaged data, rerun BSA program with all fitting parameters fixed to their bootstrap averages to reconstruct the scaling function and test the goodness of fit via $\chi^{2}_{\mathrm{red}}$.
In practice, the only user-specified settings in the above workflow are the initial guesses and their associated jitter ranges. For the correlation-ratio collapse ($Y=R$), we we initialize $U_{c,0}=3.75$ and $c_{1,0}=0.9$ for the Hubbard model, and $V_{c,0}=1.35$ and $c_{1,0}=1.1$ for the $t$-$V$ model. For the analysis of $m^{2}_{\mathrm{AFM}}$ ($G_{AB}$) versus $R$, only the exponent $c_{2}$ is jittered, with initial values $c_{2,0}=-1.7$ (0.15-0.15) for the Hubbard model and $c_{2,0}=-1.5$ (0.08-0.08) for the $t$-$V$ model. The uniform jitter width is set to 0.1 for all parameters, i.e., $\Delta U_{c}=\Delta V_{c}=\Delta c_{1}=\Delta c_{2}=0.1$. We perform 1000 bootstrap trials for each fitting window and confirm that increasing $N_{\mathrm{boot}}$ does not significantly change the results.
IV.4 Acknowledgments
We thank J. A. Gracey and T. C. Lang for helpful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the National Natural Science Foundation of China (Grants No. 12447103, No. 12274289 and No. 125B2077), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and Shanghai Jiao Tong University 2030 Initiative. F.-H. W. is supported by T.D. Lee scholarship. The computations in this paper were partially run on the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University.
Author contributions: X.Y.X. conceptualized the study and supervised the project; F.-H. W. performed the calculations and analyzed the data with the help of F. S. and C. H.; F.-H. W. and X.Y.X. wrote the manuscript.
Competing interests: The authors declare no competing interests.
Data availability: The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.